# Re3Sim: Generating High-Fidelity Simulation Data via 3D-Photorealistic Real-to-Sim for Robotic Manipulation

### v2

@@ 4.2 Rendering in Robot Task
Table 1: Background rendering methods comparison. Quantitative results of the reconstruction result in terms of PSNR and SSIM. Results are computed from 8 robot arm positions in both the wrist and third-person views.
@@
Polycam $11.57$$\pm 1.25$ $0.30$$\pm 0.08$
3DGS $\boldsymbol{14.02}$$\boldsymbol{\pm 1.50}$ $\boldsymbol{0.34}$$\boldsymbol{\pm 0.05}$
High-quality visual results not only require high-quality reconstruction but also precise alignment. To address Q1, we evaluate the quality and alignment of reconstruction results by selecting a trajectory from the simulation dataset of tasks in Sec. 4.3. This trajectory is then replayed in the real world to obtain ground truth (GT) images captured by calibrated cameras. We compare the visual similarity from both the wrist view and a third-person perspective, showcasing qualitative rendering results in Fig. 3. More rendering results in the simulation are in Sub.  E. The high similarity to real-world images further demonstrates the accuracy of our alignment. Furthermore, Tab. 1 shows the PSNR (peak signal-to-noise ratio) and SSIM (structural similarity index measure) evaluating the quantitative vision gap of the background rendering method used in Re3Sim against the reconstruction and rendering method used in Torne et al. (2024), which reconstruct the background by Polycam and render the mesh to get the visual observations. These results demonstrate the advantage of the design choice of the rendering component in Re3Sim, which benefits small vision gaps.
@@ 4.6 Pipeline Efficiency
Time consumption breakdown. We measured the average time spent on each phase during the data collection of the pick and drop a bottle task in the simulator over 60 steps, as shown in Tab. 5. Rendering 480p images from two camera views adds 12.93 milliseconds, representing one-third of the total time per step. Re3Sim operates at approximately 24 frames per second (FPS) with two cameras, ensuring real-time performance and visual fidelity.

### v3

@@ 4.2 Rendering in Robot Task
Table 1: Background rendering methods comparison. Quantitative results of the reconstruction result in terms of PSNR and SSIM. Results are computed from 1029 robot arm positions in the wrist view.
@@
Polycam $11.52$$\pm 1.40$ $0.34$$\pm 0.04$
OpenMVS $\boldsymbol{13.40}$±$\boldsymbol{\pm 0.96}$ $0.27$$\pm 0.03$
3DGS $\boldsymbol{13.29}$$\pm\boldsymbol{1.11}$ $\boldsymbol{0.37}$$\boldsymbol{\pm 0.04}$
High-quality visual results not only require high-quality reconstruction but also precise alignment. To address Q1, we evaluate the quality and alignment of reconstruction results by selecting a trajectory from the simulation dataset of tasks in Sec. 4.3. This trajectory is then replayed in the real world to obtain ground truth (GT) images captured by calibrated cameras. We compare the visual similarity from both the wrist view and a third-person perspective, showcasing qualitative rendering results in Fig. 3. More rendering results in the simulation are in Sub.  E. The high similarity to real-world images further demonstrates the accuracy of our alignment. Furthermore, Tab. 1 shows the PSNR (peak signal-to-noise ratio) and SSIM (structural similarity index measure) evaluating the quantitative vision gap. We compare Re3Sim with the reconstruction methods OpenMVS and PolyCam (the method used in  Torne et al. (2024)). Both methods reconstruct textured meshes, which can be rendered to generate visual observations. The objects are manually aligned with those in the simulation, and some pixel-level discrepancies remain. The background alignment also has some pixel-level deviations. These factors collectively lead to the relatively low PSNR and SSIM values, especially in the texture-rich scene. 3DGS outperforms Polycam in both RSNR and SSIM. Its PSNR is comparable to OpenMVS, but SSIM is notably higher. Qualitative visual comparisons and more analysis are shown in Appendix B. OpenMVS’s reconstruction has cracks, causing an obvious sim-to-real gap. These results demonstrate the advantage of the design choice of the rendering component in Re3Sim, which benefits small vision gaps.
@@ 4.6 Pipeline Efficiency
Time consumption breakdown. We measured the average time spent on each phase during the data collection of the pick and drop a bottle task in the simulator over 60 steps, as shown in Tab. 5. The reconstructed mesh, with numerous vertices and faces, demands more time for physics simulation. Rendering 480p images from two camera views using 3DGS adds 12.93 milliseconds, representing one-third of the total time per step. Re3Sim operates at approximately 24 frames per second (FPS) with two cameras, ensuring real-time performance and visual fidelity.
@@ Appendix B More Experiment Details
Figure 6: Visual comparison of rendering approaches. Rendering results of reconstruction outputs from PolyCam, OpenMVS, and 3DGS, compared with real-world photos.
Qualtative comparison of background rendering methods. We collected a trajectory using teleoperation in the simulator and replayed it in the real world. Visual comparisons of reconstruction results from various methods are shown in Figure  6. PolyCam tends to blur details in certain areas, OpenMVS’s reconstructed mesh texture shows cracks, while 3DGS produces high-quality renderings. Table  1 shows that OpenMVS’s PSNR is slightly higher than 3DGS’s, likely because the background and crack colors are similar, making PSNR less sensitive. However, the cracks cause increased variance within patches, leading to a lower SSIM compared to 3DGS.